# Text2PDE: Latent Diffusion Models for Accessible Physics Simulation

### v1

@@ E.1 Cylinder Flow
The original data are generated using COMSOL, with 600 timesteps from $t=0$ to $t=6$ on a spatial domain $(x=[0,1.6],y=[0,0.4])$, with varying inlet velocities and cylinder sizes/positions. To speed up our experiments, we downsample along the temporal dimension to 25 timesteps. We use 1000 samples for training and 100 samples for validation.
@@ Appendix G Inference Time Comparison
Model Params Inference Time (s)
GINO 72M 0.112
GNN 101M 1.683
OFormer 131M 1.336
LDMS-FF 198M 15.753
LDMM-FF 667M 27.895
LDMS-Text 313M 16.251
LDMM-Text 804M 29.096
@@
Model Params Inference Time (s)
FNO 510M 1.512
Unet 580M 2.326
Dil-Resnet 33.2M 5.827
ACDM 404M 317.872
LDMS-FF 243M 16.863
LDMM-FF 725M 25.233
LDML-FF 1.55B 57.442
LDMS-Text 334M 16.294
LDMM-Text 825M 26.494
LDML-Text 2.69B 84.457
@@
Table 8: Inference time to predict single temporal rollout (batch size = 1), evaluated on a single NVIDIA RTX 6000 Ada GPU. Reported times are averaged over nine validation samples.
@@
For transparency, we report inference times of all models on both the Cylinder Flow and Smoke Buoyancy benchmarks in Table 8. We reproduce previous results that GINO is an efficient architecture for unstructured problems (Li et al., 2023c). Furthermore, we reproduce previous results demonstrating the increased compute needed for dilated Resnets (Li et al., 2023b). Lastly, we find that our latent diffusion framework is much faster than prior work on autoregressive diffusion or learning directly in the pixel space.
@@
However, we recognize that the increased latent diffusion inference times may be a potential limitation. The current implementation uses 1000 denoising steps, which leaves ample room for accelerating inference. Indeed, the latent diffusion is on the order of 10 times slower than deterministic methods. This is a margin that using 10 times fewer denoising steps (100 steps) may be able to address, and without significant loss of accuracy if using modern techniques such as distillation. Furthermore, diffusion sampling may still be faster than using numerical counterparts.
@@
This is a challenging claim to quantify, as the numerical solver needs to be coarsened to match the accuracy of the neural solver, as well as a reasonable effort needs to be made to compare against a state of the art numerical solver (McGreivy & Hakim, 2024). The solvers used for the Cylinder Flow and Smoke Buoyancy benchmarks are COMSOL and PhiFlow, which are both high-performance numerical packages. Unfortunately, we do not have access to the COMSOL code to solve the cylinder flow problem with reduced accuracy. However, the original authors report a speedup of 11-290x (Pfaff et al., 2021), which may be enough margin to overcome coarsening the simulation to 90% of its original accuracy (see Table 1, .04 L1 Loss corresponds to about 10% normalized L1 error for this dataset).
@@
However, for the smoke buoyancy benchmark, we have access to PhiFlow and, in fact, use it to re-solve text-conditioned samples. We compare the accuracy and speed of using PhiFlow at lower resolutions with latent diffusion models in Table 9. PhiFlow was solved faster on our CPU (AMD Ryzen Threadripper PRO 5975WX 32-Cores) than on our GPU (NVIDIA RTX 6000 Ada), so numerical solver times are reported using the CPU. When running lower-resolution PhiFlow simulations, the original initial condition was downsampled, evolved with the solver, and upsampled using bicubic interpolation. There are tangible accuracy and speed gains from using a latent diffusion model; however, faster inference techniques will likely be needed to vastly outperform numerical solvers. Also of note is that text2PDE generation is something that numerical solvers cannot do.

### v2

@@ E.1 Cylinder Flow
Figure 11: Distribution of Reynolds numbers within the training set. Samples with Karman Vortex streets are also identified.
The original data are generated using COMSOL, with 600 timesteps from $t=0$ to $t=6$ on a spatial domain $(x=[0,1.6],y=[0,0.4])$, with varying inlet velocities and cylinder sizes/positions. To speed up our experiments, we downsample along the temporal dimension to 25 timesteps. We use 1000 samples for training and 100 samples for validation. Both datasets contain samples with Reynolds numbers ranging from 100-1500. Additionally, we identify Karman vortex streets in data samples by calculating the norm of the temporal velocity derivative $||\frac{dv_{x}}{dt}+\frac{dv_{y}}{dt}||_{2}$ and comparing it to a cutoff; if this quantity remains large or increases over time then periodic vortex shedding continues to occur and if this quantity decreases or goes to zero over time then the solution has reached a laminar steady state. The Reynolds number of training data samples, as well as Karman vortex street behavior is plotted in Figure 11. The observed Karman vortex streets within this dataset are consistent with the literature describing its occurrence at Reynolds numbers between 300-2e5 (Chen, 1985; Gabbai & Benaroya, 2005).
@@ E.2 Smoke Buoyancy
E.3 3D Turbulence
We use data from Lienen et al. (2024), which is generated from LES simulation in OpenFOAM with a viscosity of $\nu=10^{-5}$. The original data is given at a resolution of $192\times 48\times 48$ and at 5000 timesteps. A variety of shapes (steps, corners, pillars, teeth, bars, elbows, donut, U-shape, T-shape, H-shape, squares, crosses, platforms, etc.) are used to model flows around different objects, and for additional visualizations we refer readers to page 18 of the original paper (Lienen et al., 2024). Details on the calculation of the log-TKE distance can also be found in the original paper.
@@ F.2 Smoke Buoyancy
F.3 3D Turbulence
All baselines were trained on a single NVIDIA A100 GPU with a learning rate of $10^{-5}$ until convergence. LDM models were trained on four NVIDIA A100 80GB GPUs were used to handle memory requirements. For each model, we describe the key hyperparameters.
FNO
We train a FNO model with a hidden dimension of 192, 16 modes, and 6 layers.
Factformer
We train a Factformer model with a hidden dimension of 256, 8 heads, and 18 layers.
Dil-Resnet
We train a Resnet with a hidden dimension of 128, 4 layers, and a dilation of up to 8.
LDM
We train an autoencoder with a hidden dimension of 64 and 4 downsampling layers, resulting in a latent size of $6\times 12\times 3\times 3$, which is a compression ratio of 4096. The DiT backbone uses 1000 denoising steps, a patch size of (1, 1, 2, 2), a hidden size of 2048, and a depth of 28.
@@ Appendix G Inference Time Comparison
Model Params Time
GINO 72M 0.11
MGN 101M 1.68
OFormer 131M 1.34
LDMS-FF 198M 15.75
LDMM-FF 667M 27.89
LDMS-Text 313M 16.25
LDMM-Text 804M 29.10
@@
Model Params Time
FNO 510M 1.52
Unet 580M 2.33
Dil-Resnet 33.2M 5.83
ACDM 404M 317.8
LDMS-FF 243M 16.86
LDMM-FF 725M 25.23
LDML-FF 1.55B 57.44
LDMS-Text 334M 16.30
LDMM-Text 825M 26.49
LDML-Text 2.69B 84.46
@@
Model Params Time
FNO 1.02B 1.57
FactFormer 41.4M 5.15
Dil-Resnet 24.8M 38.8
LDML-FF 2.72B 75.2
LDML-Text 2.73B 75.7
(c) Inference Time for 3D Turbulence
@@
Table 11: Inference time to predict single temporal rollout (batch size = 1), evaluated on a single NVIDIA RTX 6000 Ada GPU or NVIDIA A100 GPU (3D Turbulence). Reported times are in seconds and are averaged over nine validation samples.
@@
For transparency, we report inference times of all models on the Cylinder Flow, Smoke Buoyancy, and 3D Turbulence benchmarks in Table 11. We reproduce previous results that GINO is an efficient architecture for unstructured problems (Li et al., 2023c). Furthermore, we reproduce previous results demonstrating the increased compute needed for dilated Resnets (Li et al., 2023b). Lastly, we find that our latent diffusion framework is much faster than prior work on autoregressive diffusion or learning directly in the pixel space.
@@
However, we recognize that the increased latent diffusion inference times may be a potential limitation. The current implementation uses 1000 denoising steps, which leaves ample room for accelerating inference. Indeed, using a DDIM sampler, as shown in Appendix C.3, is able to vastly accelerate inference and outperform conventional, deterministic neural solvers in both speed and accuracy.
@@
To compare the proposed method to a numerical baseline, the numerical solver needs to be coarsened to match the accuracy of the neural solver, as well as a reasonable effort needs to be made to compare against a state of the art numerical solver (McGreivy & Hakim, 2024). The solvers used for the Cylinder Flow, Smoke Buoyancy, and 3D Turbulence benchmarks are COMSOL, PhiFlow and OpenFOAM, which are all high-performance numerical packages. We do not have access to the COMSOL code to solve the cylinder flow problem with reduced accuracy. However, the original authors report a speedup of 11-290x (Pfaff et al., 2021), which may be enough margin to overcome coarsening the simulation to 90% of its original accuracy (see Table 1, model is reported with a 13% relative L2 error). Additionally, we also do not have access to the OpenFOAM simulation, but the original authors report a simulation time of around 10-20 minutes on 16 CPU cores (Lienen et al., 2024).
@@
For the smoke buoyancy benchmark, we have access to PhiFlow and, in fact, use it to re-solve text-conditioned samples. We compare the accuracy and speed of using PhiFlow at lower resolutions with latent diffusion models in Table 12. PhiFlow was solved faster on our CPU (AMD Ryzen Threadripper PRO 5975WX 32-Cores) than on our GPU (NVIDIA RTX 6000 Ada), so numerical solver times are reported using the CPU. When running lower-resolution PhiFlow simulations, the original initial condition was downsampled, evolved with the solver, and upsampled using bicubic interpolation. There are tangible accuracy and speed gains from using a latent diffusion model; however, faster inference techniques such as DDIM are needed to vastly outperform numerical solvers. Also of note is that text2PDE generation is something that numerical solvers cannot do.